# Modeling and Numerical Analysis of Kangaroo Lower Body based on Constrained Dynamics of Hybrid Serial-Parallel Floating-Base Systems

### 07 Dec 2023 v1

@@ 4 Analysis of Kangaroo Lower Body
The lower body of the Kangaroo robot consists of two legs, each one modeled with $32$ passive rotational DOFs and $6$ actuated linear DOFs. Therefore, the total number of DOFs in the Kangaroo model is $76$, comprising of $64$ passive DOFs and $12$ active DOFs. [^4] Here, we are not considering the floating-base DOFs. The lower body prototype weighs $\approx 42\ [kg]$ with the protection cages for the torso, not present in Fig. 4, and $\approx 36\ [kg]$ without them. All the linear actuators, except for those located at the thighs, are equipped with single-turn absolute encoders on the rotary motors allowing a relative measurement of the linear screw displacement. The thigh actuators, on the other hand, are equipped with absolute linear encoders. To calibrate the relative linear encoders, an initial calibration is performed using absolute encoders mounted at certain passive DOFs.
@@
  • 1DOF RRPR Hip Yaw;
  • 2DOFs U-2RRPU Hip Pitch/Roll;
  • 1DOF RRRR-RRRP-R Knee;
@@
We will now discuss each sub-mechanism separately, introducing a brief analysis of the kinematics singularities and the countermeasures taken to avoid them in the design of the platform. Table 2 reports an overview of each closed sub-mechanism with the number of active and passive DOFs (between parenthesis are reported the total number of passive DOFs including those that have been opened for model the closed linkages), the number of constraints to model the linkage, the actuators limits, the DOFs limits, and the average human limits at the same DOF (see brockett2016biomechanics ; lee2014design ).
Table 2: Overview of the number of parallel sub-mechanism based modules in Kangaroo lower body
@@
Ankle Roll Ankle Pitch 2 16(24) 16 -0.02 0.02 -27 -34 27 36 -12 -15 23 45
@@
Recalling Fig. 3, a first set of singularities appear when the constraint ${}^{a}\vartheta_{u}^{1}$ and ${}^{a}\vartheta_{u}^{2}$ reach the $0\ [rad]$ and $-\pi\ [rad]$ configurations, respectively when the two linear actuators are simultaneously driven to approximately $-0.07\ [m]$ and $0.05\ [m]$. Nevertheless, both singularities lie outside the linear actuator limits, $(-0.04,0.04)\ [m]$. Unfortunately, a singularity arises when the two linear actuators are driven differentially to their extremes, such as $-0.04\ [m]$ and $0.04\ [m]$, causing $\theta_{u,8}$ to approach $-\frac{\pi}{2}\ [rad]$ or $\frac{\pi}{2}\ [rad]$. To address this issue, a mechanical limit has been implemented for the Hip Roll, which can be enforced at the actuator level via the QP formulation in (4).
@@
Figure 8: 1DOF RRRR-RRRP-R Knee simplified sub-mechanisms: a. the RRRR located near the hip; b. the RRRP at the knee; c. the RRRRRR that connects the tibia to the hip. Each sub-mechanism introduces 2 linear constraints. Notice that some passive DOFs are shared between the sub-mechanisms.
@@
The Ankle consists of a hybrid serial-parallel system formed by the three sub-mechanisms depicted in Fig. 9.
Figure 9: 2DOFs U-2(RRRP-RRR-UU) Ankle Pitch/Roll simplified sub-mechanisms: a. the 2RRRP placed at the hip, that introduces $4$ constraints; b. the 2RRRR transferring the motion from the hip to the knee, that introduces again $4$ constraints; c. the U-2UR at the foot that introduces $8$ constraints. Notice that also in this complex hybrid serial-parallel chain, some passive DOFs are shared between the linkages.
@@
4.1 Non-Linear Transmission Analysis
@@
Figure 14: Passive and active DOFs during retraction and extension of the leg while keeping constant the Ankle pitch joint. The violet and yellow lines show that to preserve the joint’s position (orange line) throughout the motion, adjustments to the actuated Ankle DOFs are also required.
Finally, Fig. 15 reports the non-linear transformation between Leg Length actuator force to foot sole normal force, while performing a squat motion and applying the maximum force at the actuator ($5000\ N$). It is worth noticing that at the nominal configuration, the foot’s sole normal force is between its minimum and maximum, when completely compressed it reaches its minimum, while at half compression, it reaches its maximum. From these plots, we can see that the non-linear transmission has been designed to be able to exert its maximum force when half-squatting, for example during a jump preparation. Being torque and speed inversely related, the non-linear transmission design was optimized to achieve high force when the robot is still standing at a normal height and high speed when the leg is crouched to accelerate for jumping efficiently and maximise take off velocity.
@@
5 Comparison with TALOS
In this section, we conduct a comparison with another full-size humanoid bipedal system, the TALOS robot, developed by PAL Robotics stasse2017talos . TALOS features a more traditional serial kinematics design for its legs, encompassing 6 actuated DOFs per leg, with motors distributed along the entire chain.
@@ 5 Comparison with TALOS
with $\bm{\Lambda}\in\mathbb{R}^{6\times 6}$, $\mathbf{J}$ the Jacobian of $\mathcal{F}_{l}$ in $\mathcal{F}_{b}$, $\mathbf{M}$ the joint space inertia matrix, $\mathbf{I}\in\mathbb{R}^{6\times 6}$ identity matrix and $\epsilon$ a regularisation term used when the Jacobian is near singularities. For serial-parallel hybrid linkages, it is possible to consider the same equation (31) using the projected versions of the Jacobian and inertia matrices (see Appendix B. Projected Floating-Base Dynamics).
@@
Taking inspiration from the study conducted in Ramos22 , we extend our comparison to encompass the Centroidal Angular Momentum Matrices (CAMM) of Kangaroo and TALOS. The CAMM can be computed from the System Momentum Matrix (SMM), defined in orin2013centroidal as the product between the System Inertia and System Jacobian:
@@
with $\mathbf{A}_{S}\in\mathbb{R}^{6N\times n+6}$ the System Momentum, $N+1$ is the number of the links of the system, $\mathbf{B}_{S}\in\mathbb{R}^{6n\times 6n}$ the System Inertia matrix and $\mathbf{J}_{S}\in\mathbb{R}^{6n\times n+6}$ the System Jacobian (see orin2013centroidal for details). The SMM in (33) can be projected at the centroid of the system such that $\mathbf{A}_{G}=\mathbf{X}^{T}_{G}\mathbf{A}_{S}$, with $\mathbf{X}_{G}\in\mathbb{R}^{6N\times 6}$ the stacking of adjoint matrices transforming motions from centroidal coordinates to link coordinates, and $\mathbf{A}_{G}\in\mathbb{R}^{6\times n+6}$ defined as the Centroidal Momentum Matrix (CMM), that maps joint velocities into centroidal momentum $\mathbf{h}=\mathbf{A}_{G}\bm{\nu}\in\mathbb{R}^{6}$, with the CAMM the angular part of the CMM. As for the equivalent Cartesian Inertia matrix, it is possible to compute an equivalent CMM projected onto the closed linkage constraints.
@@
Table 3: Comparison among Kangaroo and Talos
@@
A succinct overview of the findings from this comparison is presented in Table 3. The results highlight a substantial enhancement in the Kangaroo platform in comparison to Talos. This improvement pertains to both the equivalent Cartesian inertia observed at the feet and the Centroidal Angular Momentum Matrix (CAMM).
6 Experiments using the Full-Model
This section reports preliminary experiments on the real platform employing the full-model and applying algorithms derived from the constrained modeling presented in Section 3. The Kangaroo full-model is constructed following the Unified Robotics Description Format (URDF) guidelines. However, the default URDF parser [^6] https://github.com/ros/urdfdom disallows joints with identical parent or child links, preventing the direct modeling of closed linkages. In contrast, the Simulation Description Format (SDF) [^7] http://sdformat.org/ permits such joints, enabling the simulation of closed linkages via the Gazebo Simulator koenig2004design . For this reason, we designed a configuration YAML file and a novel parser that permits the use of the URDF of Kangaroo with control libraries (e.g. the CartesI/O framework, by Laurenzi19 ) where all the necessary closed-linkage constraints are listed. Finally, we developed a C++ library for the kinematics and dynamics modeling of series-parallel hybrid chains named Closed Linkage Library (CLL) based on RBDL felis2017rbdl and Eigen eigenweb . The following experiments have been carried out using the CLL library on an Intel Core i7 CPU 2.30GHz × 16 cores with 32 GiB RAM, and are presented in the video accompanying this paper.
6.1 Computation of Passive Kinematics Quantities and Initial Calibration
In general, when modeling serial-parallel hybrid linkages as constrained multi-body systems, it is necessary to have knowledge of the complete state vector denoted by $\bm{\theta}$ and $\bm{\dot{\theta}}$. However, this information may not be entirely accessible due to some passive quantities being either fully or partially unmeasurable. On the other hand, the actuated quantities typically remain observable. Furthermore, the linear actuators in Kangaroo are based on ball-screws, equipped with single turn absolute encoders at the motor side, permitting a relative measurement of the linear displacement of the screw, hence working as linear relative encoders. For this reason, it is necessary to properly offset the initial value of the linear actuators to have a linear absolute reading, at every robot initialization. This calibration is made according to the arrangement of the closed kinematics and the measurements from the absolute encoders mounted on some of the passive DOFs.
@@ 6.1 Computation of Passive Kinematics Quantities and Initial Calibration
6.2 Quasi-Static Contact Wrench Estimation
@@
where we have used (3.1). Reorganizing leads to:
@@
$=\mathbf{J}_{c,m}(\mathbf{q})^{T}\mathbf{F},$ (6.2)
@@
with $\mathbf{J}_{c,m}(\mathbf{q})\in\mathbb{R}^{3c\times a}$ and $(\cdot)^{\dagger}$ a properly computed pseudo-inverse. Notice that velocity measurements could be included to consider the effect of Coriolis/centrifugal terms. From the wrench estimation, we compute the Zero Moment Point (ZMP) being a fundamental quantity used in bipedal locomotion for stabilization, see Alex12 ; kajita2014introduction .
@@
Each linear actuator is equipped with a load cell to measure the ball-screw axial force $\bar{\tau}$. Furthermore, the axial force is also estimated from the motor torque $\tau_{m}$:
@@
with $\eta=0.95$ the efficiency of the ball screw provided by the constructor, and $L$ the screw lead, that for the leg length actuator is $0.1\ [m]$ and for the hip and ankle actuators is $0.05\ [m]$. The motor torque, in turn, is computed from the motor current. Fig. 21(a) reports the reconstructed contact wrench filtered using a low-pass second-order Butterworth filter with cuttoff frequency $5\ [Hz]$. It is possible to see that the estimated tangential forces correctly sum up at the weight of the robot without the cage, approximately $36\ [kg]$. In Fig. 21(b) are reported the computed ZMP and feet Centers of Pressure (COPs) using the estimated contact wrenches. The contact wrench algorithms estimation time is $\approx 0.03\ [ms]$, using the fullPivLu [^9] https://eigen.tuxfamily.org/dox/classEigen_1_1FullPivLU.html method from the Eigen library, and dense matrices representation, allowing for fast real-time implementation.
@@
7 Conclusion and Future Works
@@
Future works will regard the deployment of Whole-Body Inverse Dynamics on the Kangaroo platform together with model predictive control algorithms to perform highly agile and dynamic motions, such as running and jumping. In particular, we envision the employment of a simplified, serial model of the robot for the model predictive control, i.e. considering only significative passive DOFs, and mapping the control input onto the full model using the IK and ID presented in this paper. We are also investigating the extension of passivity-based approaches, see cic , for Operational Space Control to series-parallel hybrid chains. Finally, the tools developed in this work will be the basis for the design and development of the upper body for the Kangaroo robot, to evaluate different kinematic structures for the torso and the arms.
Appendix
A. Computation of Relative Kinematics
@@
A1. Relative Velocity and Jacobian
@@
$\begin{bmatrix}{{}^{w}\mathbf{v}_{a,b}}\\ {{}^{w}\bm{\omega}_{a,b}}\end{bmatrix}=\begin{bmatrix}{{}^{a}\mathbf{R}_{w}}&\mathbf{0}\\ \mathbf{0}&{{}^{a}\mathbf{R}_{w}}\end{bmatrix}\begin{bmatrix}{{}^{w}\mathbf{v}_{b}}\\ {{}^{w}\bm{\omega}_{b}}\end{bmatrix}-\mathbf{A}\begin{bmatrix}{{}^{w}\mathbf{v}_{a}}\\ {{}^{w}\bm{\omega}_{a}}\end{bmatrix},$ (48)
@@
${{}^{w}\mathbf{v}_{a,b}}$ $={{}^{a}\mathbf{R}_{w}}{{}^{w}\mathbf{v}_{b}}-{{}^{a}\mathbf{R}_{w}}{{}^{w}\mathbf{v}_{a}}+{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\mathbf{p}_{a,b}}\times{{}^{w}\bm{\omega}_{a}}\right)=$
@@
${{}^{w}\bm{\omega}_{a,b}}$ $={{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\omega}_{b}}-{{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\omega}_{a}}={{}^{a}\mathbf{R}_{w}}\left({{}^{w}\bm{\omega}_{b}}-{{}^{w}\bm{\omega}_{a}}\right)=$
@@
A2. Relative Acceleration and Acceleration Bias

### 22 Feb 2024 v2

@@ IV Analysis of Kangaroo Lower Body
The lower body of the Kangaroo robot consists of two legs, each one modeled with $32$ passive rotational DOFs and $6$ actuated linear DOFs. Therefore, the total number of DOFs in the Kangaroo model is $76$, comprising of $64$ passive DOFs and $12$ active DOFs. [^4] Here, we are not considering the floating-base DOFs. The lower body prototype weighs $\approx 42\ [kg]$ with the protection cages for the torso, not present in Fig. 4, and $\approx 36\ [kg]$ without them. All the linear actuators, except for those located at the thighs, are equipped with relative linear encoders, while the thigh actuators are equipped with absolute linear encoders. To calibrate the relative linear encoders, an initial calibration is performed using absolute encoders mounted at certain passive DOFs.
@@
  • 1DOF RRPR Hip Yaw,
  • 2DOFs U-2RRPU Hip Pitch/Roll,
  • 1DOF RRRR-RRRP-R Knee,
@@
We will now discuss each sub-mechanism separately, introducing a brief analysis of the kinematics singularities and the countermeasures taken to avoid them in the design of the platform. Table II reports an overview of the closed sub-mechanisms with the number of active and passive DOFs (with the number of total passive DOFs), the number of constraints to model the linkage, the actuators limits, the DOFs limits, and the average human limits at the same DOF (see [42, 43]).
TABLE II: Overview of the number of parallel sub-mechanism based modules in Kangaroo lower body
@@
Ankle Roll Ankle Pitch 2 16(24) 16 -0.02 0.02 -27 -36 27 34 -45 -12 15 23
@@
Recalling Fig. 3, a first set of singularities appear when the constraint ${}^{a}\vartheta_{u}^{1}$ and ${}^{a}\vartheta_{u}^{2}$ reach the $0\ [rad]$ and $-\pi\ [rad]$ configurations, respectively when the two linear actuators are simultaneously driven to approximately $-0.07\ [m]$ and $0.05\ [m]$. Nevertheless, both singularities lie outside the linear actuator limits, $(-0.04,0.04)\ [m]$. Unfortunately, a singularity arises when the two linear actuators are driven differentially to their extremes, such as $-0.04\ [m]$ and $0.04\ [m]$, causing $\theta_{u,8}$ to approach $-\frac{\pi}{2}\ [rad]$ or $\frac{\pi}{2}\ [rad]$. To address this issue, a physical constraint has been implemented for the Hip Roll, which can be enforced at the actuator level via the QP formulation in (IV).
@@
Figure 8: 1DOF RRRR-RRRP-R Knee simplified sub-mechanisms: a. RRRR introducing $2$ constraints; b. RRRP, introducing $2$ constraints; c. RRRRRR introducing $2$ constraints. Some passive DOFs are shared between the linkages.
@@
The Ankle consists of a hybrid serial-parallel sub-mechanism formed by the three sub-mechanisms depicted in Fig. 9.
Figure 9: 2DOFs U-2(RRRP-RRR-UU) Ankle Pitch/Roll simplified sub-mechanisms: a. 2RRRP placed at the hip, introducing $4$ constraints; b. 2RRRR transferring the motion from the hip to the knee introducing $4$ constraints; c. U-2UR introducing $8$ constraints. Some passive DOFs are shared between the linkages.
@@
IV-A Non-Linear Transmission Analysis
@@
Figure 14: Passive and active DOFs during retraction and extension of the leg while keeping constant the Knee pitch joint. The violet and yellow lines show that to preserve the joint’s position (orange line) throughout the motion, adjustments to the actuated Ankle DOFs are also required.
Finally, Fig. 15 reports the non-linear transformation between Leg Length actuator force to foot sole normal force, while performing a squat motion and applying the maximum force at the actuator ($5000\ N$). It is worth noticing that at the nominal configuration, the foot’s sole normal force is between its minimum and maximum, when completely compressed it reaches its minimum, while at half compression, it reaches its maximum. From these plots, we can see that the non-linear transmission has been designed to be able to exert its maximum force when half-squatting, for example during a jump preparation. Being torque and speed inversely related, the non-linear transmission design was optimized to achieve high force when the robot is still standing at a normal height and high speed when the leg is crouched to accelerate for jumping efficiently.
@@
V Comparison with TALOS
In this section, we conduct a comparison with another full-size humanoid bipedal system, the TALOS robot, developed by PAL Robotics [12]. TALOS features a more traditional serial kinematics design for its legs, encompassing 6 actuated DOFs per leg, with motors distributed along the entire chain.
@@ V Comparison with TALOS
with $\bm{\Lambda}\in\mathbb{R}^{6\times 6}$, $\mathbf{J}$ the Jacobian of $\mathcal{F}_{l}$ in $\mathcal{F}_{b}$, $\mathbf{M}$ the joint space inertia matrix, $\mathbf{I}\in\mathbb{R}^{6\times 6}$ identity matrix and $\epsilon$ a regularisation term used when the Jacobian is near singularities. For serial-parallel hybrid linkages, it is possible to consider the same equation (31) using the projected versions of the Jacobian and inertia matrices (see Appendix B).
@@
Taking inspiration from the study conducted in [44], we extend our comparison to encompass the Centroidal Angular Momentum Matrices (CAMM) of Kangaroo and TALOS. The CAMM can be computed from the System Momentum Matrix (SMM), defined in [45] as the product between the System Inertia and System Jacobian:
@@
with $\mathbf{A}_{S}\in\mathbb{R}^{6N\times n+6}$ the System Momentum, $N+1$ is the number of the links of the system, $\mathbf{B}_{S}\in\mathbb{R}^{6n\times 6n}$ the System Inertia matrix and $\mathbf{J}_{S}\in\mathbb{R}^{6n\times n+6}$ the System Jacobian (see [45] for details). The SMM in (33) can be projected at the centroid of the system such that $\mathbf{A}_{G}=\mathbf{X}^{T}_{G}\mathbf{A}_{S}$, with $\mathbf{X}_{G}\in\mathbb{R}^{6N\times 6}$ the stacking of adjoint matrices transforming motions from centroidal coordinates to link coordinates, and $\mathbf{A}_{G}\in\mathbb{R}^{6\times n+6}$ defined as the Centroidal Momentum Matrix (CMM), that maps joint velocities into centroidal momentum $\mathbf{h}=\mathbf{A}_{G}\bm{\nu}\in\mathbb{R}^{6}$, with the CAMM the angular part of the CMM. As for the equivalent Cartesian Inertia matrix, it is possible to compute an equivalent CMM projected onto the closed linkage constraints.
@@
TABLE III: Comparison among Kangaroo and Talos
@@
A succinct overview of the findings from this comparison is presented in Table III. The results highlight a substantial enhancement in the Kangaroo platform in comparison to Talos. This improvement pertains to both the equivalent Cartesian inertia observed at the feet and the Centroidal Angular Momentum Matrix (CAMM).
VI Preliminary Experiments with Full-Model
This section reports preliminary experiments on the real platform employing the full-model and applying algorithms derived from the constrained modeling presented in Section III. The Kangaroo full-model is constructed following the Unified Robotics Description Format (URDF) guidelines. However, the default URDF parser [^6] https://github.com/ros/urdfdom disallows joints with identical parent or child links, preventing the direct modeling of closed linkages. In contrast, the Simulation Description Format (SDF) [^7] http://sdformat.org/ permits such joints, enabling the simulation of closed linkages via the Gazebo Simulator [46]. For this reason, we designed a configuration YAML file and a novel parser that permits the use of the URDF of Kangaroo with control libraries (e.g. the CartesI/O framework, by [47]) where all the necessary closed-linkage constraints are listed. Finally, we developed a C++ library for the kinematics and dynamics modeling of series-parallel hybrid chains named Closed Linkage Library (CLL) based on RBDL [48] and Eigen [49]. The following experiments have been carried out using the CLL library on an Intel Core i7 CPU 2.30GHz × 16 cores with 32 GiB RAM, and are presented in the video accompanying this paper.
VI-A Computation of Passive Kinematics Quantities and Initial Calibration
In general, when modeling serial-parallel hybrid linkages as constrained multi-body systems, it is necessary to have knowledge of the complete state vector denoted by $\bm{\theta}$ and $\bm{\dot{\theta}}$. However, this information may not be entirely accessible due to some passive quantities being either fully or partially unmeasurable. On the other hand, the actuated quantities typically remain observable. Furthermore, the linear actuators in Kangaroo are based on ball-screws, equipped with absolute encoders at the motor side, permitting a relative measurement of the linear displacement of the screw, hence working as linear relative encoders. For this reason, it is necessary to properly offset the initial value of the linear actuators to have a linear absolute reading, at every robot initialization. This calibration is made according to the arrangement of the closed kinematics and the measurements from the absolute encoders mounted on some of the passive DOFs.
@@ VI-A Computation of Passive Kinematics Quantities and Initial Calibration
VI-B Quasi-Static Contact Wrench Estimation
@@
where we have used (III-A). Reorganizing leads to:
@@
$=\mathbf{J}_{c,m}(\mathbf{q})^{T}\mathbf{F},$ (VI-B)
@@
with $\mathbf{J}_{c,m}(\mathbf{q})\in\mathbb{R}^{3c\times a}$ and $(\cdot)^{\dagger}$ a properly computed pseudo-inverse. Notice that velocity measurements could be included to consider the effect of Coriolis/centrifugal terms. From the wrench estimation, we compute the Zero Moment Point (ZMP) being a fundamental quantity used in bipedal locomotion for stabilization, see [50, 51].
@@
The estimation of the axial force $\bar{\tau}$ in each linear actuator is computed from the measured motor torque $\tau_{m}$ as:
@@
with $\eta=0.95$ the efficiency of the ball screw provided by the constructor, and $L$ the screw lead, that for the leg length actuator is $0.1\ [m]$ and for the hip and ankle actuators is $0.05\ [m]$.
@@
Fig. 21(a) reports the reconstructed contact wrench filtered using a low-pass second-order Butterworth filter with cuttoff frequency $5\ [Hz]$. It is possible to see that the estimated tangential forces correctly sum up at the weight of the robot without the cage, approximately $36\ [kg]$. In Fig. 21(b) are reported the computed ZMP and feet Centers of Pressure (COPs) using the estimated contact wrenches. The contact wrench algorithms estimation time is $\approx 0.03\ [ms]$, using the fullPivLu [^9] https://eigen.tuxfamily.org/dox/classEigen_1_1FullPivLU.html method from the Eigen library, and dense matrices representation, allowing for fast real-time implementation.
VII Conclusion and Future Works
@@
Future works will regard the deployment of Whole-Body Inverse Dynamics on the Kangaroo platform together with model predictive control algorithms to perform highly agile and dynamic motions, such as running and jumping. In particular, we envision the employment of a simplified, serial model of the robot for the model predictive control, i.e. considering only significative passive DOFs, and mapping the control input onto the full model using the IK and ID presented in this paper. We are also investigating the extension of passivity-based approaches, see [52], for Operational Space Control to series-parallel hybrid chains. Finally, the tools developed in this work will be the basis for the design and development of the upper body for the Kangaroo robot, to evaluate different kinematic structures for the torso and the arms.
Appendix A Computation of Relative Kinematics
@@
A-A Relative Velocity and Jacobian
@@
$\begin{bmatrix}{{}^{a}\mathbf{v}_{a,b}}\\ {{}^{a}\bm{\omega}_{a,b}}\end{bmatrix}=\begin{bmatrix}{{}^{a}\mathbf{R}_{w}}&\mathbf{0}\\ \mathbf{0}&{{}^{a}\mathbf{R}_{w}}\end{bmatrix}\begin{bmatrix}{{}^{w}\mathbf{v}_{b}}\\ {{}^{w}\bm{\omega}_{b}}\end{bmatrix}-\mathbf{A}\begin{bmatrix}{{}^{w}\mathbf{v}_{a}}\\ {{}^{w}\bm{\omega}_{a}}\end{bmatrix},$ (48)
@@
${{}^{a}\mathbf{v}_{a,b}}$ $={{}^{a}\mathbf{R}_{w}}{{}^{w}\mathbf{v}_{b}}-{{}^{a}\mathbf{R}_{w}}{{}^{w}\mathbf{v}_{a}}+{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\mathbf{p}_{a,b}}\times{{}^{w}\bm{\omega}_{a}}\right)=$
@@
${{}^{a}\bm{\omega}_{a,b}}$ $={{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\omega}_{b}}-{{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\omega}_{a}}={{}^{a}\mathbf{R}_{w}}\left({{}^{w}\bm{\omega}_{b}}-{{}^{w}\bm{\omega}_{a}}\right)=$
@@
A-B Relative Acceleration and Acceleration Bias